# Diffusion Models for Intelligent Transportation Systems: A Survey

### v1

@@ Abstract
Intelligent Transportation Systems (ITS) are vital in modern traffic management and optimization, significantly enhancing traffic efficiency and safety. Recently, diffusion models have emerged as transformative tools for addressing complex challenges within ITS. In this paper, we present a comprehensive survey of diffusion models for ITS, covering both theoretical and practical aspects. First, we introduce the theoretical foundations of diffusion models and their key variants, including conditional diffusion models and latent diffusion models, highlighting their suitability for modeling complex, multi-modal traffic data and enabling controllable generation. Second, we outline the primary challenges in ITS and the corresponding advantages of diffusion models, providing readers with a deeper understanding of the intersection between ITS and diffusion models. Third, we offer a multi-perspective investigation of current applications of diffusion models in ITS domains, including autonomous driving, traffic simulation, trajectory prediction, and traffic safety. Finally, we discuss state-of-the-art diffusion model techniques and highlight key ITS research directions that warrant further investigation. Through this structured overview, we aim to provide researchers with a comprehensive understanding of diffusion models for ITS, thereby advancing their future applications in the transportation domain.
@@ I Introduction
In the past few decades, researchers have employed various approaches to address the challenges of ITS. For example, Recurrent Neural Networks (RNNs) are often used to model temporal relationships, while Convolutional neural networks (CNNs) are commonly utilized to capture spatial structure [2]. And graph-based approaches have demonstrated superior capabilities in extracting spatial correlations within traffic networks [3, 4]. However, these approaches often exhibit limitations when handling noisy or incomplete data. In contrast, generative models such as Generative Adversarial Networks (GANs) and Variational Autoencoders (VAEs) have proven effective for traffic data generation and imputation tasks [5, 6]. However, GANs suffer from unstable training, and VAE has the limitation of low-quality output. Diffusion models, as a powerful class of generative models, offer advantages such as ease of training, enhanced generative performance, controllable generation, and multi-modal capabilities. To date, diffusion models have been applied across a wide range of vision tasks [7], with promising applications such as Sora [8]. Inspired by these developments, an increasing number of researchers in the ITS domain have begun to adopt diffusion models to address various challenges in ITS. Therefore, originating in image processing and computer vision, diffusion models are now being applied across various traffic tasks, from autonomous driving and traffic simulation to traffic forecasting and traffic safety. As illustrated in Fig. 1, diffusion models are suitable for processing various traffic data, and are capable of addressing a wide range of various traffic tasks based on task-specific conditions or unconditional methods.
@@
There have been numerous surveys on ITS [2, 9], as well as specific technologies within the ITS domain [3, 4, 10]. Similarly, several reviews have focused on diffusion models [11, 12, 13] and their applications in areas such as computer vision [7] and medical imaging [14]. However, there is currently no comprehensive review of diffusion models within the ITS domain.
@@
To address this gap, this paper presents a detailed literature review on diffusion models in ITS. First, we outline how diffusion models have emerged as powerful tools for various traffic tasks. Specifically, we introduce the theoretical foundations of diffusion models, along with conditional diffusion models and latent diffusion models, which extend their applicability to more specific tasks within ITS. Second, we examine the key challenges in ITS and the corresponding advantages of diffusion models. Third, we investigate the applications of diffusion models in areas such as autonomous driving, traffic simulation, traffic forecasting, and traffic safety within ITS, as shown in Fig. 6. In particular, we review these applications based on criteria such as task, denoising condition, or model architecture, as illustrated in Table. I. Finally, we provide an outlook on potential future directions for diffusion models in ITS. Our goal is to bridge the gap between the diffusion model and transportation research communities, fostering interdisciplinary collaboration and advancing the application of diffusion models in transportation.
@@
  • We systematically introduce how diffusion models have become powerful approaches for various traffic tasks by processing multi-modal and complex traffic data. Additionally, we explore the key challenges in ITS and the corresponding advantages of diffusion models. This analysis offers readers deeper insights into the intersection of ITS and diffusion models.
@@
  • We discuss the cutting-edge techniques in diffusion models, and highlight key research directions for diffusion models in ITS that are worthy of further exploration.
The remainder of the paper is organized as follows: Sec. II presents theoretical foundations of diffusion models and their key variants. Sec. III outline the key challenges in ITS and the corresponding advantages of diffusion models. Sec. IV- Sec. VII explores the diverse applications of diffusion models within ITS, including autonomous driving, traffic simulation, traffic forecasting, and traffic safety. Sec. VIII discusses several promising directions for future research. Finally, the conclusions are drawn in Sec. IX.
@@ II THEORY
Diffusion models have emerged as transformative tools in the field of ITS. This section outlines how diffusion models have become powerful and flexible methods for addressing various traffic-related challenges. First, we explore the theoretical foundations of diffusion models, which lies in their ability to learn the underlying data distribution through a process of noise injection and subsequent denoising. This makes them highly effective for modeling complex traffic dynamics. Next, we introduce key variants of diffusion models, particularly conditional diffusion models and latent diffusion models, which extend their applicability to more specific and challenging tasks within ITS. By incorporating domain-specific conditions and leveraging latent spaces, diffusion models can be applied to multi-modal traffic data, offering solutions to a wide range of traffic-related tasks.
@@ II-A Foundations of Diffusion Models
Diffusion models are a powerful class of probabilistic generative models that gradually perturb data by adding Guassian noise to data, and then learn to reverse this process to generate new data. During training, the model learns to denoise the data at each step, effectively transforming random noise into coherent and realistic outputs.
This section provides an overview of three predominant formulations in diffusion models: Denoising Diffusion Probabilistic Models (DDPMs), which utilize discrete steps to incrementally add and remove noise; Noise Conditioned Score Networks (NCSNs), which estimate the gradient of the log-density of the data distribution to guide sample generation; and Stochastic Differential Equations (SDEs), which offer a continuous-time perspective that unifies and generalizes both DDPMs and NCSNs under a common mathematical framework.
@@ II-A1 Denoising Diffusion Probabilistic Models (DDPMs)
DDPMs [15, 16] utilize two Markov chains: a forward (diffusion) process that gradually adds Gaussian noise to data, transforming it into pure noise over multiple steps, and a reverse (denoising) process, learned through neural networks—typically based on a U-Net architecture [17]—that progressively removes the noise to reconstruct the original data.
@@
Reverse (Denoising) Process. The goal of DDPMs is to learn the reverse of this diffusion process, where the model starts with Gaussian noise and progressively removes the noise to generate new data. The reverse process is also modeled as a Markov chain, but it is parameterized by a neural network $p_{\theta}({x}_{t-1}|{x}_{t})$ that generates $p_{\theta}({x}_{0})$ in a step-by-step manner:
@@ II-A2 Noise Conditioned Score Networks (NCSNs)
NCSNs [18] are a class of score-based generative models that focus on estimating the score function of the data distribution. Instead of explicitly modeling the reverse diffusion process, NCSNs learn the gradient of the log-density of the data distribution at various noise levels via score matching [19], and subsequently generate samples via Langevin dynamics [20, 21].
Score Matching. Given an unknown data distribution $p_{data}(x)$, the score function of the data density $p_{(}x)$ is defined as $\nabla_{x}\log p(x)$. The score network $\mathbf{s}_{\theta}$, a neural network parameterized by $\theta$, is trained to estimate the score function $\nabla_{x}\log p(x)$. When the data distribution is unknown, score estimation can be performed using sliced score matching [22] or denoising score matching [23]. In NCSNs [18], denoising score matching is adopted, wherein data are perturbed with multiple levels of Gaussian noise. Specifically, the noise distribution is pre-specified as $q_{\sigma}(\tilde{x}|x)=\mathcal{N}(\tilde{x}|x,\sigma^{2}\mathbf{I})$, and the gradient of the log-likelihood with respect to the noisy data is given by $\nabla_{\tilde{x}}\log q_{\sigma}(\tilde{x}|x)=-(\tilde{x}-x)/\sigma^{2}$. Given a sequence noise scales $\sigma_{1}<\sigma_{2}<...<\sigma_{L}$, the denoising score matching objective for all $\sigma\in\{\sigma_{i}\}^{L}_{i=1}$ is defined as:
@@
Langevin Dynamics. To generate samples, NCSNs employ annealed Langevin dynamics, starting with large noise levels and gradually annealing down to lower noise levels. At each noise level, Langevin dynamics is iteratively applied using the learned score function to recover the original data distribution progressively. The update rule for Langevin dynamics is given by:
@@
Figure 2: Different condition mechanisms for diffusion models. (1) Concatenation-based mechanism incorporates conditions such as historical data and maps directly into the input. (2) Cross-attention-based mechanism integrates conditions like text and external features through cross-attention layers. (3) Classifier-based mechanism uses an external classifier to guide denoising based on conditions such as reinforcement learning or cost functions. (4) Classifier-free mechanism combines conditional and unconditional denoising models, balancing both with a weight parameter.
@@ II-B Variants of Diffusion Models
In this section, we introduce key variants of diffusion models, including conditional diffusion models and latent diffusion models (LDMs), which have significantly advanced the field of intelligent transportation systems. These models not only enhance the ability to generate realistic traffic data but also offer flexibility and controllability in modeling complex traffic environments. By incorporating domain-specific information, such as historical data, traffic layouts, or external semantic features, conditional diffusion models enable the generation of more accurate and diverse traffic scenarios that reflect real-world conditions. Meanwhile, LDMs operate in a lower-dimensional latent space, facilitating faster training and inference times while maintaining the fidelity of generated outputs. Additionally, LDMs allows multi-modal conditions within the latent space. These capabilities make LDMs particularly useful for image-based, video-based, or text-involved traffic tasks. Collectively, these advanced models demonstrate the potential of diffusion models to revolutionize intelligent transportation systems, providing powerful tools for traffic simulating, forecasting, and optimization in increasingly dynamic urban environments.
@@ II-B1 Conditional Diffusion Models
The three types of standard diffusion models introduced above are unconditional, where the inputs are limited to the perturbed data $x_{t}$ and the diffusion step $t$. Conditional diffusion models, on the other hand, incorporate conditional information as an extra input, allowing for control over the generation process according to specific requirements. This capability makes them highly adaptable for various applications in intelligent transportation systems. Below, we focus on four primary conditioning mechanisms: concatenation-based, cross-attention-based, classifier-based, and classifier-free-based approaches. Concatenation-based methods are simple to implement but may struggle to capture complex relationships between the data and conditions. Cross-attention-based methods excel at modeling long-range dependencies and complex interactions, and enable multi-modal conditioning, but they do not offer control over the strength of the conditions. Classifier-based approaches provide adjustable guidance through external classifiers but can be limited by the accuracy and generalization capability of the classifier. Classifier-free-based methods are flexible and do not require additional classifiers, but they often come with increased training costs. The visualization of these four conditioning mechanisms is shown in Fig. 2.
Concatenation-based. In concatenation-based mechanisms, the conditioning information is directly concatenated with the perturbed data $x_{t}$ or the diffusion step $t$, and then fed into the model for sample generation. This method is simple and effective, allowing the model to leverage the conditioning information throughout the denoising process. For example, in the field of intelligent transportation systems, conditioning on historical data [27, 28, 29] or map feature [30] has been employed for generating traffic trajectories. Similarly, image features [31, 32] or traffic layout [33] have been directly concatenated with the noise data vector for generating traffic scenarios. Additionally, conditioning on trip regions [34], road network [35], or graph structure [36] has been applied in traffic flow generation. These examples emphasize the effectiveness of concatenation-based mechanisms in diverse transportation applications.
@@
Cross-attention-based. Cross-attention-based conditional diffusion models integrate the cross-attention layers [37] into the denoising networks, enabling effective fusion of conditioning information during the denoising process and guiding the network to generate outputs aligned with the conditions. The cross-attention mechanism plays an important role in facilitating the interaction between the conditioning information and the noisy data, especially in scenarios where their relationship is complex or involves different modalities, such as text and images. Stable Diffusion [38] introduced a general-purpose conditioning mechanism based on cross-attention, enabling multi-modal conditional inputs, making diffusion models into powerful and flexible generators. Building on this foundational work, numerous studies have applied this cross-attention-based conditioning mechanism in the field of intelligent transportation systems. For example, conditioning on text [39, 40, 32, 33, 41, 42], drive actions [30], external features and semantic features [43], origin-destination-departure time (ODT) feature [44], or bounding boxes [42] has been used for various traffic-related tasks.
@@
Following this work, many studies related to traffic trajectory generation and motion planning have designed various classifiers to controllably generate traffic scenarios that comply with traffic rules and ensure trajectory smoothness. For example, the cumulative rewards learned through reinforcement learning [46], motion planning cost function [47], STL formulas based on traffic rules [48], language-based loss function [49], and driving behavior classes [50] have been designed as classifier to generate task-conditioned samples.
@@
For many traffic-related generation tasks, researchers have employed the classifier-free guidance mechanism to regulate the diversity of the generated outputs [52, 50, 53, 54, 55, 34]. This approach prevents the outputs from following the conditional guidance too closely or being constrained too tightly.

### v3

@@ Abstract
Intelligent Transportation Systems (ITS) play a crucial role in enhancing traffic efficiency and safety. Recently, diffusion models have emerged as transformative tools for addressing the complex challenges faced within ITS. This paper presents a comprehensive survey of diffusion models in ITS, exploring both theoretical and practical dimensions. We begin by introducing the theoretical foundations of diffusion models and their key variants, such as conditional and latent diffusion models, emphasizing their capacity to model intricate, multi-modal traffic data and enable controllable generation. Next, we outline the primary challenges in ITS and the advantages diffusion models provide, facilitating a deeper understanding of the intersection between diffusion models and ITS. We then conduct a multi-perspective examination of current applications of diffusion models across ITS domains, including autonomous driving, traffic simulation, traffic forecasting, and traffic safety. Finally, we discuss state-of-the-art diffusion model techniques and highlight key research directions within ITS that merit further exploration. Through this structured overview, we aim to equip researchers with a comprehensive understanding of diffusion models in ITS, thereby fostering their future applications in the transportation domain.
@@ I Introduction
In the past few decades, researchers have employed various approaches to address the challenges of ITS. For example, Recurrent Neural Networks (RNNs) are often used to model temporal relationships, while Convolutional neural networks (CNNs) are commonly utilized to capture spatial structure [2]. And graph-based approaches have demonstrated superior capabilities in extracting spatial correlations within traffic networks [3, 4]. However, these approaches often exhibit limitations when handling noisy or incomplete data. In contrast, generative models such as Generative Adversarial Networks (GANs) and Variational Autoencoders (VAEs) have proven effective for traffic data generation and imputation tasks [5, 6]. However, GANs suffer from unstable training, and VAE has the limitation of low-quality output. As a powerful class of generative models, diffusion models offer advantages such as ease of training, enhanced generative performance, controllable generation, and multi-modal capabilities. To date, diffusion models have been applied across a wide range of vision tasks [7], with promising applications such as Sora [8]. Inspired by these developments, an increasing number of researchers in the ITS domain have begun to adopt diffusion models to address various challenges in ITS. Therefore, originating in image processing and computer vision, diffusion models are now being applied across various traffic tasks, including autonomous driving, traffic simulation, traffic forecasting, and traffic safety. As illustrated in Fig. 1, diffusion models are suitable for processing various traffic data and can address a wide range of traffic tasks based on task-specific conditions or unconditional methods.
@@
There have been numerous surveys on intelligent transportation systems (ITS), focusing on deep learning techniques [2, 9], along with focused surveys on graph-based learning methods [3, 4] and generative AI approaches [10] for ITS applications. Meanwhile, several reviews have focused on diffusion models [11, 12, 13] and their applications in areas such as computer vision [7] and medical imaging [14]. However, there is currently no comprehensive review of diffusion models within the ITS domain.
@@
To address this gap, this paper presents a detailed literature review on diffusion models in ITS. First, we outline how diffusion models have emerged as powerful tools for various traffic tasks. Specifically, we introduce the theoretical foundations of diffusion models, along with conditional diffusion models and latent diffusion models, which extend their applicability to more specific tasks within ITS. Second, we examine the critical challenges in ITS and the corresponding advantages of diffusion models. Third, we investigate the applications of diffusion models in areas such as autonomous driving, traffic simulation, traffic forecasting, and traffic safety within ITS, as shown in Fig. 6. In particular, we review these applications based on criteria such as task, denoising condition, or model architecture, as illustrated in Table. I. Finally, we provide an outlook on potential future directions for diffusion models in ITS. Our goal is to bridge the gap between the diffusion model and transportation research communities, fostering interdisciplinary collaboration and advancing the application of diffusion models in transportation.
@@
  • We systematically introduce how diffusion models have become powerful approaches for various traffic tasks by processing multi-modal and complex traffic data. Additionally, we explore the critical challenges in ITS and the corresponding advantages of diffusion models. This analysis offers readers more profound insights into the intersection of ITS and diffusion models.
@@
  • We discuss the cutting-edge techniques in diffusion models and highlight key research directions for diffusion models in ITS that are worthy of further exploration.
The remainder of the paper is organized as follows: Sec. II presents theoretical foundations of diffusion models and their key variants. Sec. III outline the key challenges in ITS and the corresponding advantages of diffusion models. Sec. IV-A- Sec. IV-D explores the diverse applications of diffusion models within ITS, including autonomous driving, traffic simulation, traffic forecasting, and traffic safety. Sec. V discusses several promising directions for future research. Finally, the conclusions are drawn in Sec. VI.
@@ II THEORY
Diffusion models have emerged as transformative tools in the field of ITS. This section outlines how diffusion models have become powerful and flexible methods for addressing various traffic-related challenges. First, we explore the theoretical foundations of diffusion models, which lie in their ability to learn the underlying data distribution through a process of noise injection and subsequent denoising. This makes them highly effective for modeling complex traffic dynamics. Next, we introduce key variants of diffusion models, particularly conditional and latent diffusion models, which extend their applicability to more specific and challenging tasks within ITS. By incorporating domain-specific conditions and leveraging latent spaces, diffusion models can be applied to multi-modal traffic data, offering solutions to a wide range of traffic-related tasks.
@@ II-A Foundations of Diffusion Models
Diffusion models are a powerful class of probabilistic generative models that gradually perturb data by adding Gaussian noise to data and then learn to reverse this process to generate new data. During training, the model learns to denoise the data at each step, effectively transforming random noise into coherent and realistic outputs.
This section provides an overview of three predominant formulations in diffusion models: Denoising Diffusion Probabilistic Models (DDPMs), which utilize discrete steps to add and remove noise incrementally; Noise Conditioned Score Networks (NCSNs), which estimate the gradient of the log-density of the data distribution to guide sample generation; and Stochastic Differential Equations (SDEs), which offer a continuous-time perspective that unifies and generalizes both DDPMs and NCSNs under a common mathematical framework.
@@ II-A1 Denoising Diffusion Probabilistic Models (DDPMs)
DDPMs [15, 16] utilize two Markov chains: a forward (diffusion) process that gradually adds Gaussian noise to data, transforming it into pure noise over multiple steps, and a reverse (denoising) process, learning through neural networks—typically based on a U-Net architecture [17]—that progressively removes the noise to reconstruct the original data.
@@
Reverse (Denoising) Process. DDPMs aim to learn the reverse of this diffusion process, where the model starts with Gaussian noise and progressively removes the noise to generate new data. The reverse process is also modeled as a Markov chain, but it is parameterized by a neural network $p_{\theta}({x}_{t-1}|{x}_{t})$ that generates $p_{\theta}({x}_{0})$ in a step-by-step manner:
@@
In the context of ITS, DDPM-based models have been employed for multiple traffic tasks such as trajectory prediction, traffic scenario generation, and spatio-temporal data completion. The simplicity of their training objective and compatibility with structured conditioning makes them a strong foundation for generative modeling in transportation domains.
@@ II-A2 Noise Conditioned Score Networks (NCSNs)
NCSNs [18] are a class of score-based generative models that estimate the data distribution’s score function. Instead of explicitly modeling the reverse diffusion process, NCSNs learn the gradient of the log-density of the data distribution at various noise levels via score matching [19], and subsequently generate samples via Langevin dynamics [20].
Score Matching. Given an unknown data distribution $p_{data}(x)$, the score function of the data density $p(x)$ is defined as $\nabla_{x}\log p(x)$. The score network $\mathbf{s}_{\theta}$, a neural network parameterized by $\theta$, is trained to estimate the score function $\nabla_{x}\log p(x)$. When the data distribution is unknown, score estimation can be performed using sliced score matching [21] or denoising score matching [22]. In NCSNs [18], denoising score matching is adopted, wherein data are perturbed with multiple levels of Gaussian noise. Specifically, the noise distribution is pre-specified as $q_{\sigma}(\tilde{x}|x)=\mathcal{N}(\tilde{x}|x,\sigma^{2}\mathbf{I})$, and the gradient of the log-likelihood with respect to the noisy data is given by $\nabla_{\tilde{x}}\log q_{\sigma}(\tilde{x}|x)=-(\tilde{x}-x)/\sigma^{2}$. Given a sequence noise scales $\sigma_{1}<\sigma_{2}<...<\sigma_{L}$, the denoising score matching objective for all $\sigma\in\{\sigma_{i}\}^{L}_{i=1}$ is defined as:
@@
Langevin Dynamics. To generate samples, NCSNs employ annealed Langevin dynamics, starting with large noise levels and gradually annealing down to lower noise levels. At each noise level, Langevin dynamics is iteratively applied using the learned score function to progressively recover the original data distribution. The update rule for Langevin dynamics is given by:
@@
While NCSNs are less commonly used in traffic applications compared to DDPMs, their formulation provides theoretical insights into score-based generation and has inspired hybrid models for spatial-temporal graph generation in traffic flow analysis [23].
Figure 2: Different condition mechanisms for diffusion models. (1) Concatenation-based mechanism directly incorporates conditions such as historical data and maps into the input. (2) Cross-attention-based mechanism integrates conditions like text and external features through cross-attention layers. (3) Classifier-based mechanism uses an external classifier to guide denoising based on conditions such as reinforcement learning or cost functions. (4) Classifier-free mechanism combines conditional and unconditional denoising models, balancing both with a weight parameter.
@@ II-A3 Stochastic Differential Equations (SDEs)
The SDE-based formulation also provides a flexible framework for modeling traffic-related processes in continuous time, which is particularly advantageous in tasks involving irregularly sampled time series. For example, the CSDI model [27] applies conditional score-based diffusion to time series imputation, interpolation, and forecasting, and achieves state-of-the-art performance on datasets with high missingness and non-uniform sampling intervals.
@@ II-B Variants of Diffusion Models
In this section, we introduce key variants of diffusion models, including conditional diffusion models and latent diffusion models (LDMs), which have significantly advanced the field of intelligent transportation systems. These models enhance the ability to generate realistic traffic data and offer flexibility and controllability in modeling complex traffic environments. By incorporating domain-specific information, such as historical data, traffic layouts, or external semantic features, conditional diffusion models enable the generation of more accurate and diverse traffic scenarios that reflect real-world conditions. Meanwhile, LDMs operate in a lower-dimensional latent space, facilitating faster training and inference times while maintaining the fidelity of generated outputs. Additionally, LDMs allow multi-modal conditions within the latent space. These capabilities make LDMs particularly useful for image-based, video-based, or text-involved traffic tasks. These advanced models demonstrate the potential of diffusion models to revolutionize intelligent transportation systems, providing powerful tools for traffic simulating, forecasting, and optimization in increasingly dynamic urban environments.
@@ II-B1 Conditional Diffusion Models
The three types of standard diffusion models introduced above are unconditional, where the inputs are limited to the perturbed data $x_{t}$ and the diffusion step $t$. Conditional diffusion models, on the other hand, incorporate conditional information as an extra input, allowing for control over the generation process according to specific requirements. This capability makes them highly adaptable for various applications in intelligent transportation systems. Below, we focus on four primary conditioning mechanisms: concatenation-based, cross-attention-based, classifier-based, and classifier-free-based approaches. Concatenation-based methods are simple to implement but may struggle to capture complex relationships between the data and conditions. Cross-attention-based methods excel at modeling long-range dependencies and complex interactions with multi-modal conditioning, but they do not offer control over the strength of the conditions. Classifier-based approaches provide adjustable guidance through external classifiers but can be limited by the accuracy and generalization capability of the classifier. Classifier-free-based methods are flexible and do not require additional classifiers, but they often come with increased training costs. The visualization of these four conditioning mechanisms is shown in Fig. 2.
Concatenation-based. In concatenation-based mechanisms, the conditioning information is directly concatenated with the perturbed data $x_{t}$ or the diffusion step $t$, and then fed into the model for sample generation. This simple and effective method allows the model to leverage the conditioning information throughout the denoising process. This approach is effective for ITS tasks like trajectory and traffic flow prediction, where historical data are concatenated with input for more accurate modeling [28, 29, 30].
@@
Cross-attention-based. Cross-attention-based conditional diffusion models integrate the cross-attention layers [31] into the denoising networks, enabling effective fusion of conditioning information during the denoising process and guiding the network to generate outputs aligned with the conditions. The cross-attention mechanism plays an important role in facilitating the interaction between the conditioning information and the noisy data, especially in scenarios where their relationship is complex or involves different modalities, such as text and images. Stable Diffusion [32] introduced a general-purpose conditioning mechanism based on cross-attention, enabling multi-modal conditional inputs, making diffusion models into powerful and flexible generators. In ITS scenarios, cross-attention is commonly used when the condition (e.g., textual instruction) differs in modality from the input. For instance, Panacea [33] applies cross-attention to align BEV features with textual prompts for scene generation.
@@
This approach is commonly used in ITS tasks requiring controllable outputs, such as trajectory generation and traffic scenario simulation tasks, ensuring conformity to traffic rules and behavior constraints [35, 36].
@@
This method is applied in ITS tasks like traffic scenario generation, where the goal is to maintain realistic outputs while allowing for diverse possibilities [38].
Types of Conditions in ITS Applications. In intelligent transportation systems (ITS), diffusion models utilize a variety of conditions, which play critical roles in incorporating context, enforcing constraints, and improving overall model performance. We summarize the main types as follows:
  • Sensor-based features: Inputs from LiDAR or cameras [39, 33] provide rich environmental context, improving the accuracy of perception tasks such as segmentation and detection.
  • Spatial features: Information such as road networks [40], BEV features [41], or traffic layouts [42] encode spatial structure and constraints, enhancing realism and consistency in tasks like trajectory prediction and scenario simulation.
  • Temporal inputs: Historical trajectories [28, 29], departure times [43], and observation intervals provide motion context and support accurate future prediction and planning under uncertainty.
  • Semantic or symbolic conditions: Text prompts [44, 45], driving intents [46], traffic rules encoded via STL formulas [36], and behavior classes [47] enable high-level control, allowing the model to generate customized, diverse, or rule-compliant outputs.
These diverse conditions enhance the controllability, contextual awareness, and robustness of diffusion models across various ITS tasks, enhancing effectiveness across a wide range of ITS applications.